OpenStudy (anonymous):

If the sphere shown above has a radius of 4 units, then what is the approximate volume of the sphere? A. 267.95 cubic units B. 334.93 cubic units C. 100.48 cubic units D. 21.33 cubic units

OpenStudy (anonymous):

@mathstudent55

OpenStudy (lord2clash):

okay so the Formula is v=4/3*pi*Radius squared so v=3/4*3.14*4^2

OpenStudy (anonymous):

but its cubed

OpenStudy (lord2clash):

i meant cubed sorry

OpenStudy (lord2clash):

okay so the Formula is v=4/3*pi*Radius cubed so v=3/4*3.14*4^3

OpenStudy (anonymous):

but how do i put that in my calculator i only have squared

OpenStudy (lord2clash):

do you have a ^ sign on your calculator?

OpenStudy (anonymous):

no I have y*x

OpenStudy (readergirl12):

If you have a scientific calculator it would look like dis  lol

OpenStudy (lord2clash):

well i believe your computer has a calculator does it not?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

It's very easy to obtain wrong answers on a calculator unless you know and follow "Order of Operations" rules. If you have v=4/3*pi*Radius cubed a better way of writing that, for maximum clarity, would be \[V=(4/3)\pi*r^3\]

OpenStudy (mathmale):

or\[\frac{ 4 }{ 3 }\pi*r^3\]

OpenStudy (anonymous):

thanks

OpenStudy (mathmale):

First, divide 3 into 4 AND PRESS "ENTER" This forces your calculator to calculate 4/3 and to hold it in memory.

OpenStudy (mathmale):

Next, multiply your result by pi.

OpenStudy (anonymous):

can you hlp with more plz

OpenStudy (anonymous):

George has two cans containing biscuit dough. The cans are cylindrical and their dimensions (in inches) are as follows. Can A: h = 10 and r = 0.8 Can B: h = 5 and r = 1.6 Which of the following statements about the volumes of the cans is true? A. The volume of can B is four times the volume of can A B. The volume of can B is half the volume of can A C. The volumes of the two cans are equal D. The volume of can A is half the volume of can B

OpenStudy (mathmale):

Yes, but I need to excuse myself for a few minutes first.

OpenStudy (anonymous):

ok

OpenStudy (readergirl12):

Good job everyone. :D

OpenStudy (mathmale):

If you have more questions, post them, one at a time, as new posts.